OpenStudy (anonymous):

Runner A is initially 5.0 km west of a flagpole and is running with a constant velocity of 4.6 km/h due east. Runner B is initially 8.8 km east of the flagpole and is running with a constant velocity of 7.0 km/h due west. What will be the distance of the two runners from the flagpole when their paths cross?

OpenStudy (anonymous):

Convert units to m/s. \[5_{km}=5000_m\] \[8.8_{km}=8800_m\] \[\frac{ 4.6 }{ 3.6 } = 1.277_{m/s}\] \[\frac{ 7 }{ 3.6 }=1.944_{m/s}\]

OpenStudy (anonymous):